Calculate the electric field 0.245 cm from a point charge of 25.6 nC.


1
Expert's answer
2021-12-20T10:27:30-0500

The electric field of a point charge is given as follows:


"E = k\dfrac{q}{r^2}"

where "k = 9\times 10^9N\cdot m^2/C^2", "q = 25.6\times10^{-9}C", and "r = 0.245cm = 0.00245m". Thus, obtain:


"E = 9\times 10^9\cdot \dfrac{25.6\times 10^{-9}}{0.00245^2} \approx 3.84\times 10^7V/m"

Answer. "3.84\times 10^7V/m".